OpenStudy (anonymous):

The equation of circle having a diameter with endpoints (-3, 1) and (5, 7) is

OpenStudy (cwrw238):

midpoint of line joining these points is center of circle half length is radius then substitute in the general form (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

A.(x - 4)² + (y - 3)² = 25 B.(x - 1)² + (y - 4)² = 25 C.(x - 1)² + (y - 4)² = 100

OpenStudy (zzr0ck3r):

sqrt( (7-1)^2 + (5- (-3))^2 ) is the radius so (x+3)^2 + (y-1)^2 = 100

OpenStudy (anonymous):

c?

OpenStudy (zzr0ck3r):

hmm did i do something wrong? or is the answer not listed?

OpenStudy (campbell_st):

the centre is the mid point of the interval the radius is the distance from the centre to an endpoint centre: x = (5 -3)/2 y = (7 + 1)/2 centre is (1, 4) Radius: \[r = \sqrt{(5 - 1)^2 +(7 - 4)^2}\] \[r = \sqrt{18}\] then the equation is \[(x - 1)^2 + (y - 4)^2 = 18\] you can expand and simplify the equation

OpenStudy (campbell_st):

oops radius = 5 so r^2 = 25

OpenStudy (zzr0ck3r):

ahh yes "the diameter has endpoint" sorry

OpenStudy (paxpolaris):

yes ... diameter is 10 ... so radius is 5

OpenStudy (anonymous):

whats the answer then